# A Multivariate Unimodality Test Harnessing the Dip Statistic of Mahalanobis Distances Over Random Projections

### v3

@@ 4 Experiments
In this section, we present a comprehensive suite of experiments conducted for both multidimensional unimodality testing and estimation of the number of clusters in clustering problems. The decision to evaluate our algorithm for cluster estimation stems from its complexity and numerous practical applications (Schubert, 2023). A pertinent query pertains to which $\alpha$-unimodality family we aim to detect. Despite its rigor, we opted to assess 1-unimodality regardless of the underlying data space. Empirical results indicated its efficacy even on challenging real-world datasets. As previously highlighted, our algorithm is characterized by three parameters: $M,\epsilon,p$. Following an initial exploration, we identified parameter values that consistently produced favorable outcomes. Specifically, we set $M=100$, $\epsilon=0.99$, $p=0.99$, and chose a significance level $a=0.01$.
@@ 4.1 Unimodality Experiments
Table 1: Performance comparison of dip-dist (DD), mudpod (MP), and folding (F) tests in determining unimodality or multimodality. The table displays the percentage of multimodality cases identified over 10 runs. In synthetic experiments, 1000 points were sampled, akin to the MNIST experiments where 1000 points were sampled from the training set for each instance. For space constraints, Single MNIST experiment results are compressed, with digits divided by semi-colons summarizing the outcomes across all tests. $G$ denotes a 2D or 3D Gaussian distribution. $C(r)$ symbolizes the 2D equation of a circle with radius $r$. The parametric equations $U(\theta)=\left(\cos(\theta),\sin(\theta)\right)$ and $L(\theta)=\left(1-\cos(\theta),1-\sin(\theta)-0.5\right)$ with $\theta\in[0,\pi]$ are also utilized.
Table 1 presents an intricate assessment of the capability of three distinct tests — dip-dist (DD), mudpod (MP), and folding (F) — in discerning unimodal and multimodal datasets. DD and F tests are non-parametric and we also set $a=0.01$. The evaluation was carried out over ten distinct runs for each test on a combination of both synthetic and real-world data drawn from the MNIST dataset (LeCun et al., 1998). Starting with synthetic unimodal datasets, namely the single 2D and 3D Gaussian distributions, we find a unanimous agreement across the three tests, with none indicating any instances of multimodality. Examining the synthetic bimodal distributions, 2D Moons and Circles show clear multimodality, confirmed by DD and MP tests with 100% detection. DD and MP also report 100% detection for bimodal Gaussians in 2D and 3D. However, DD’s performance declines with three closely aligned Gaussians in both 2D and 3D, unlike MP’s consistent 100% detection. The F test only identifies multimodality in the 2D Circles dataset.
@@
Table 2: Ablation study examining the influence of space, distance, and observer selection strategy on the unimodality detection performance of mudpod. The table showcases the agreement of mudpod’s results for a mixture of two 2D Gaussians comprising two terms in 2D, where ground truth unimodality is well established. In the notation used, O denotes Original, RP represents Randomly Projected, E stands for Euclidean, M signifies Mahalanobis, while R and P represent Random and Percentile, respectively.
Ablation Study: Our test comprises several components, including random projections, Mahalanobis distance, and uniform sampling of distances from origin percentiles. In Table 2, we present an ablation study to evaluate the significance of each component. We have generated datasets from a mixture of two 2D Gaussians as the target distribution, for which the bimodality or unimodality can be determined analytically (Konstantellos, 1980). The reported performance is aggregated from a series of experiments conducted with four different significance levels $0.001,0.005,0.01,0.05$, across 1000 distinct data sets, with each experiment executed 10 times. Our ablation study reveals key insights into the performance of different observer picking strategies, spaces, and distances for unimodality detection. Primarily, the percentile strategy (P) consistently for observer picking demonstrated superiority over the random strategy (R) across all tested combinations of space and distance. Furthermore, the Mahalanobis distance consistently emerged as a more effective metric compared to the Euclidean distance. This performance difference was especially evident in the Randomly Projected space, suggesting that the intrinsic characteristics of the Mahalanobis distance, e.g., accounting for data covariance, plays a pivotal role in enhancing detection reliability. Notably, the randomly projected (RP) space exhibited a pronounced advantage over the original space in our assessments. This superiority held true irrespective of the distance metric or observer strategy employed. Such a trend strongly indicates that the RP space aligns more coherently with the ground truth unimodality, offering a more reliable landscape for detection when contrasted with the original space. In summary, our findings recommend a strategic combination of the P strategy, Mahalanobis distance, and the RP space.
@@
Table 3: The table presents the number of clusters ($K$) and the associated Normalized Mutual Information (NMI) values obtained by various methods on different datasets. Values marked with † could not be computed due to memory constraints or were terminated after 8 hours. All results are represented as mean ± standard deviation over 10 executions. It is noteworthy that for the k-means algorithm, the correct number of clusters was always predefined.
Figure 1: The plot illustrates the relative error between the estimated and actual values of the number of clusters $K$ generated by mp-means, as the number of Monte Carlo simulations increases. Each experiment was conducted over 10 executions, with the variance also depicted. The plot is more discernible in color.
@@ 4.2 Clustering Experiments
This section analyzes the performance of mp-means, an approach that incorporates mud-pod into the dip-means wrapper method, replacing dip-dist. This wrapper procedure begins with a single cluster and iteratively splits those clusters deemed by a multivariate unimodality test as likely having multiple subclusters. The procedure terminates when no further splits are suggested. A similar integration of the folding test yields the fold-means algorithm. In Table 3, we compare the performance of various clustering methods estimating the number of clusters across several datasets. It is important to note that in all our experiments, we use the raw flattened data encoding and apply only a feature-wise z-transformation. Although our method employs the Mahalanobis distance and is scale-agnostic, we observed that scaling profoundly affects other algorithms. Our experiments include datasets like USPS, MNIST, Fashion-MNIST (F-MNIST), Human Activity Recognition (HAR), Optdigits, Pendigits, Isolet, and TCGA. Detailed dataset descriptions are provided in Appendix A. As ground truth number of clusters, the number of classes was considered in all datasets. We benchmark against classic algorithms, including x-means (Pelleg & Moore, 2000), g-means (Pelleg & Moore, 2000), pg-means (Feng & Hamerly, 2006), dip-means (Kalogeratos & Likas, 2012), and hdbscan (Campello et al., 2013). A detailed overview of the algorithms and their hyperparameter setups are provided in Appendix B. Our evaluation metrics consist of the estimated number of clusters (k) and the Normalized Mutual Information (NMI). NMI values lie between [0, 1], where 1 signifies a perfect match and 0 represents an arbitrary result.
@@
This experiment set, encompassing both the complexity of many modes and inherent clustering errors, is notably more challenging than single-class unimodality testing. We observe that most of the classical approaches, notably x-means and g-means, have the tendency to predict a high value for $k$, frequently estimating around 35 clusters for datasets such as USPS, MNIST, and F-MNIST. Similarly, x-means, for instance, consistently overestimates the $k$ value on these datasets, accompanied by NMI values which, while respectable, do not always achieve the top tier. In contrast, pg-means often significantly underestimates the number of clusters, suggesting a notable disparity in its clustering from the ground truth. This assertion is further supported by its suboptimal NMI values across multiple datasets, with values as low as 0.14 for the USPS dataset. Hdbscan demonstrates a varying range in $k$ values across datasets, highlighting its adaptability. However, this variability does not always correlate with high NMI values. Its performance on the USPS dataset, where it estimates 13 clusters with an NMI of 0.38, is a case in point.
@@
Turning our attention to the HAR dataset, the dip-means and mp-means methods are particularly notable. They estimate $k$ values closely aligned with the ground truth of 5 and simultaneously achieve the highest NMI scores, specifically 0.73. This is indicative of their robustness and accuracy in deciphering the cluster structure inherent to the data. In the context of the Optdigits and Pendigits datasets, dip-means struggles to align its $k$ estimations with the ground truth, predicting values of 1 and 16, respectively. Conversely, mp-means appears more consistent, with $k$ values of 8 and 14 for Optdigits and Pendigits, respectively, and corresponding NMI scores that are commendable. In an experiment with k-means using accurate cluster count, the NMI of mp-means closely aligns with, or even exceeds, that of k-means. Finally, the projection dimension of mp-means varies between subclusters. The experiments underscore its robustness across multiple projection spaces. In conclusion, while various algorithms demonstrate strengths in certain areas, the consistent performance of mp-means across multiple datasets underscores its utility in solving clustering problems with unknown number of clusters.
@@
Ablation Study on Simulations’ Number: In Figure 1, the relative error between the estimated number of clusters $K$ and the true value, as deduced by the mp-means algorithm, is depicted against the number of Monte Carlo simulations ($M$). The shaded regions around each line represent the variance over 10 executions, highlighting result consistency. As observed, for all datasets, an increase in the number of Monte Carlo simulations tends to correspond with a decline in the relative error, signifying an enhancement in the accuracy of $k$ estimation. While some datasets exhibit minimal spread, others display significant variance, suggesting the need for more Monte Carlo simulations. This suggests a need for more projections on certain datasets. However, as noted in the table, this variance does not significantly impede $k$ estimation. Summarizing, Figure 1 substantiates the notion that increasing the number of Monte Carlo simulations augments the precision of the mp-means algorithm’s cluster estimation, albeit with varying degrees of improvement across different datasets.
@@ 5 Conclusions
In this study, we addressed the challenges associated with generalizing the concept of unimodality testing to higher dimensions. Through the innovative integration of random projections and the notion of $\alpha$-unimodality, we have presented a novel methodology which offers a robust approach to multivariate unimodality testing. Leveraging both point-to-point distancing and linear random projections, our approach bridges the gap between the simplicity of one-dimensional unimodality confirmation and the intricacies of its higher-dimensional counterpart. By integrating our unimodality test with k-means, we introduced a new incremental clustering approach that automatically estimates the cluster count while performing clustering. Empirical evaluations underscored the efficacy of our test in both unimodality assessments and clustering scenarios, showcasing its indispensable applicability in diverse data-driven domains. Future exploration will focus to additional $\alpha$-unimodality preserving operations to enhance our test, experimenting with $\alpha$-unimodality for $\alpha\neq 1$, and applying the proposed test to other data analysis tasks, e.g., capturing covariate shift, time series change detection, etc.

### v4

@@ 4 Experiments
In this section, we present a comprehensive suite of experiments conducted for both multivariate unimodality testing and estimating cluster counts in clustering tasks. Our decision to assess our algorithm for cluster estimation is driven by its complexity and wide practical relevance (Schubert, 2023). A pertinent query pertains to which $\alpha$-unimodality family we aim to detect. Despite its rigor, we opted to assess 1-unimodality regardless of the underlying data space. Empirical results indicated its efficacy even on challenging real-world datasets. Our algorithm is characterized by three parameters: $M,\epsilon,p$. Following an initial exploration, we identified parameter values that consistently produced favorable outcomes. Specifically, we set $M=100$, $\epsilon=0.99$, $p=0.99$, and chose a significance level $a=0.01$.
@@ 4.1 Unimodality Experiments
Table 1: Performance comparison of dip-dist (DD), mudpod (MP), and folding (F) tests in determining unimodality or multimodality. The table displays the percentage of multimodality cases identified over 10 runs. 1000 points randomly sampled from synthetic sets and MNIST training set per experiment. For space constraints, Single MNIST experiment results are compressed, with digits divided by semi-colons summarizing the outcomes across all tests. $G_{n}$ denotes a $n$D Gaussian distribution. $C(r)$ symbolizes the 2D equation of a circle with radius $r$. Utilizing parametric equations $U(\theta)=\left(\cos(\theta),\sin(\theta)\right)$ and $L(\theta)=\left(1-\cos(\theta),1-\sin(\theta)-0.5\right)$ with $\theta\in[0,\pi]$.
@@
Table 1 presents an intricate assessment of the capability of three distinct tests — dip-dist (DD), mudpod (MP), and folding (F) — in discerning unimodal and multimodal datasets. DD and F tests are non-parametric and we also set $a=0.01$. It is important to note that DD is a special case of mud-pod, omitting $\alpha$ exponent, operating in the original space using the Euclidean distance and considering all data points as observers. The evaluation was carried out over ten distinct runs for each test on a combination of both synthetic and real-world data drawn from the MNIST dataset (LeCun et al., 1998). Starting with synthetic unimodal datasets, namely the single 2D and 3D Gaussian distributions, we find a unanimous agreement across the three tests, with none indicating any instances of multimodality. Examining the synthetic bimodal distributions, 2D Moons and Circles show clear multimodality, confirmed by DD and MP tests with 100% detection. DD and MP also report 100% detection for bimodal Gaussians in 2D and 3D. However, DD’s performance declines with three closely aligned Gaussians in both 2D and 3D, unlike MP’s consistent 100% detection. The F test only identifies multimodality in the 2D Circles dataset.
@@
Table 2: Impact of space, distance, and observer selection strategy on mudpod’s unimodality detection performance. Mudpod’s result agreement is shown for a mixture of two 2D Gaussians with confirmed ground truth unimodality. Notation: ‘O’ for Original, ‘RP’ for Randomly Projected, ‘E’ for Euclidean, ‘M’ for Mahalanobis; ‘R’ and ‘P’ denote Random and Percentile, respectively.
@@
Ablation Study: Our test comprises several components, including random projections, Mahalanobis distance, and uniform sampling of distances from origin percentiles. In Table 2, we present an ablation study to evaluate the significance of each component. We have generated datasets from a mixture of two 2D Gaussians as the target distribution, for which the bimodality or unimodality can be determined analytically (Konstantellos, 1980). The reported performance is aggregated from a series of experiments conducted with four different significance levels $0.001,0.005,0.01,0.05$, across 1000 distinct data sets, with each experiment executed 10 times. Our ablation study reveals key insights into the performance of different observer picking strategies, spaces, and distances for unimodality detection. Primarily, the percentile strategy (P) for observer picking demonstrated superiority over the random strategy (R) across all tested combinations of space and distance. Furthermore, the Mahalanobis distance consistently emerged as a more effective metric compared to the Euclidean distance. This performance difference was especially evident in the Randomly Projected space, suggesting that the intrinsic characteristics of the Mahalanobis distance, e.g., accounting for data covariance, plays a pivotal role in enhancing detection reliability. Notably, the randomly projected (RP) space exhibited a pronounced advantage over the original space in our assessments. This superiority held true irrespective of the distance metric or observer strategy employed. Such a trend strongly indicates that the RP space aligns more coherently with the ground truth unimodality, offering better detection capabilities compared to the original space. In summary, our findings recommend a strategic combination of the P strategy, Mahalanobis distance, and the RP space.
Table 3: The table presents the number of clusters ( K ) and the associated NMI values obtained by various methods on different datasets. Values marked with † could not be computed due to memory constraints or were terminated after 8 hours. All results are represented as mean ± standard deviation over 10 executions. For the k-means algorithm, the correct number of clusters was always predefined.
@@
Figure 1: The plot shows the relative error between the estimated and actual number of clusters generated by mp-means, against increasing Monte Carlo simulations. Ten executions per experiment, variance depicted. The plot is more discernible in color.
@@ 4.2 Clustering Experiments
This section analyzes the performance of mp-means, an approach that incorporates mud-pod into the dip-means wrapper method, replacing dip-dist. Specifically, both mp-means and dip-means employ incremental $k$-means clustering based on testing clusters for unimodality. Starting with one cluster (the entire dataset), they incrementally increase $k$ by splitting multimodal clusters, terminating when all clusters are deemed unimodal. They differ in the cluster unimodality assessment. Dip-means uses the dip-test criterion, while mp-means uses our proposed unimodality test. Upon detecting a multimodal cluster, they split the one with the highest dip value. It’s split into two using the 2-means algorithm or assigning clusters at mean $\pm$ standard deviation, using the cluster’s mean and standard deviation. In our work, we opt for the latter for computational efficiency. In this way, the number of clusters is increased to $k+1$ and the $k+1$ centers are updated via $k$-means. A similar integration of the folding test yields the fold-means algorithm.
@@
In Table 3, we compare the performance of various clustering methods estimating the number of clusters across several datasets. It is important to note that in all our experiments, we use the raw flattened data encoding and apply only a feature-wise z-transformation. Although our method employs the Mahalanobis distance and is scale-agnostic, we observed that scaling profoundly affects other algorithms. Our experiments include datasets like USPS, MNIST, Fashion-MNIST (F-MNIST), Human Activity Recognition (HAR), Optdigits, Pendigits, Isolet, and TCGA. Detailed dataset descriptions are provided in Appendix A. As ground truth number of clusters, the number of classes was considered in all datasets. We benchmark against classic algorithms, including x-means (Pelleg & Moore, 2000), g-means (Pelleg & Moore, 2000), pg-means (Feng & Hamerly, 2006), dip-means (Kalogeratos & Likas, 2012), and hdbscan (Campello et al., 2013). A detailed overview of the baseline algorithms and their hyperparameter setups are provided in Appendix B. Our evaluation metrics consist of the estimated number of clusters (k) and the Normalized Mutual Information (NMI). NMI values lie between [0, 1], where 1 signifies a perfect match and 0 represents an arbitrary result.
@@
This experiment set, encompassing both the complexity of many modes and inherent clustering errors, is notably more challenging than unimodality testing. We observe that most of the classical approaches, notably x-means and g-means, tend to predict a high value for $k$, frequently estimating around 35 clusters for datasets such as USPS, MNIST, and F-MNIST. Similarly, x-means, consistently overestimates the $k$ value on these datasets, accompanied by NMI values not consistently reaching the top tier. In contrast, pg-means often significantly underestimates the number of clusters, suggesting a notable diversion from the ground truth. This assertion is further supported by its suboptimal NMI values across multiple datasets, with values as low as 0.14 for the USPS dataset. Hdbscan shows a varying range in $k$ values across datasets, highlighting its adaptability. However, this variability does not always correlate with high NMI values. Its performance on the USPS dataset, where it estimates 13 clusters with an NMI of 0.38, is a case in point.
@@
Turning our attention to the HAR dataset, the dip-means and mp-means methods are particularly notable. They estimate $k$ values closely aligned with the ground truth of 5 and simultaneously achieve the highest NMI scores, specifically 0.73. This is indicative of their capability in deciphering the cluster structure inherent to the data. In the context of the Optdigits and Pendigits datasets, dip-means struggles to align its $k$ estimations with the ground truth, predicting values of 1 and 16, respectively. Notably, mp-means appears more consistent, with $k$ values of 8 and 14 for Optdigits and Pendigits, respectively, and corresponding NMI scores that are commendable. In an experiment with k-means using accurate cluster count, the NMI of mp-means closely aligns with, or even exceeds, that of k-means. Intriguingly, the projection dimension of mp-means varies between subclusters. The experiments also underscore its robustness across multiple projection spaces. In conclusion, the consistent performance of mp-means underscores its utility in solving clustering problems with unknown number of clusters.
@@
Ablation Study on Simulations’ Number: In Figure 1, the relative error between the estimated number of clusters $K$ and the true value, as deduced by the mp-means algorithm, is depicted against the number of Monte Carlo simulations ($M$). The shaded regions around each line represent the variance over 10 executions, highlighting result consistency. As observed, for all datasets, an increase in the number of Monte Carlo simulations tends to correspond with a decline in the relative error, signifying an enhancement in the accuracy of $k$ estimation. While some datasets exhibit minimal spread, others display significant variance, suggesting the need for more Monte Carlo simulations. This suggests a need for more projections on certain datasets. However, this variance does not significantly impede $k$ estimation. Summarizing, Figure 1 substantiates the notion that increasing the number of Monte Carlo simulations augments the precision of the mp-means algorithm’s k estimation, albeit with varying degrees of improvement across different datasets. Interestingly, we note that convergence occurs at around 100 simulations, which is significant for practical use.
@@ 5 Conclusions
In this paper, we addressed the challenges associated with generalizing unimodality testing to higher dimensions. Building upon the notion of $\alpha$-unimodality, we presented a novel methodology that provides a robust approach to multivariate unimodality testing. Utilizing point-to-point distancing and linear random projections, our approach bridges the gap between the simplicity of one-dimensional unimodality confirmation and the intricacies of its higher-dimensional counterpart. By integrating our unimodality test with k-means, we introduced a new incremental clustering approach that automatically estimates the cluster count while performing clustering. Empirical evaluations affirm our test’s efficacy in unimodality assessments and clustering scenarios, highlighting its vital applicability across diverse data-driven domains. Future exploration will focus on identifying additional $\alpha$-unimodality preserving operations to enhance our test, e.g, exploiting JL lemma’s relaxations through matrix sketching, experimenting with $\alpha$-unimodality for $\alpha\neq 1$, and applying the test to various data analysis tasks, such as covariate shift and time series change detection.
@@ Appendix B Clustering Algorithms
Appendix C Implementation Details
The code implementation, developed in Python, utilizes the diptest library [^4] https://pypi.org/project/diptest/ for computing dip values and statistics. The code is publicly available at: https://github.com/prokolyvakis/mudpod.